OpenStudy (anonymous):

1) Identify each equation below as true, false, or open. a) 2x – 4 = 3x + 7 b) 4 – 7 = 7 – 4 c) –2 + 3 = 3 – 2 2) What is the solution to 6 – 5x = –9? 3) Which ordered pair is a solution of the equation y = 9 – x? a) (–6, 12) b) (–3, 12) c) (3, –6) d) (6, –3) 4) Which ordered pair is a solution of the equation y = –6x? a) (–2, 12) b) (–3, –18) c) (4, –2) d) (4, 24) 5) Which ordered pair is a solution of the equation y = 3x – 2? a) (1, –1) b) (2, 3) c) (3, 4) d) (4, 10)

OpenStudy (blank ):

1. a) open b) false c) true

OpenStudy (blank ):

2. x=3

OpenStudy (blank ):

3. (–3, 12)

OpenStudy (blank ):

4. (–2, 12)

OpenStudy (blank ):

5. (4, 10)

OpenStudy (anonymous):

6)Natasha and Alex go jogging together every Saturday morning. Natasha always runs two more miles than Alex. If the variable x represents the distance Natasha runs and y represents the distance Alex runs, graph the relationship between the distances Natasha and Alex could run each Saturday. 7) Which equation below is an identity? a) 5f + 6 = 5f – 6 b) 4q + 3 = 3q – 3 + q c) 6t + 1 – 3t = –2t – 1 d) 8a – (a + 4) = 9a – 2(a + 2) 8) Which equation below has no solution? a) 4z – 1 = –4z + 10 b) 7n + 9 = 7n – 2 c) 6s + 5 = 5s – 6 d) 8c – 3 = 8c + 7 – c 9) Chris and Lucy went to a baseball game. They each bought a bag of peanuts and Lucy purchased a bottled water. Chris had a $10 gift card toward the cost and paid the rest, which came to $32.00. If a baseball ticket costs $14.75 and a bag of peanuts costs $4.50, what is the cost of a bottled water? On problems 10-21, solve the equation. 10) –11 = q + 8 11) d – 6 = –13 12) 12.9 = 3m

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Just need help with 6-12